a company is considering building a bridge across a river. the bridge would cost $1.5 million to build and nothing to maintain. the following table shows the company's anticipated demand over the lifetime of the bridge:

price per cross: $8,$7,$6,$5,$4,$3,$2,$1,$0

number of crossing in thousand: 0,100,200,300,400,500,600,700,800

a.if the company were to build the bridge, what would be its profit-maximizing price? would that be the efficient level of output? why or why not?

b.if the company is interested in maximizing profit, should it build the bridge? what would be its profit or loss?

c.if the government were to build the bridge, what price should it charge?

d.should be government build the bridge? explain


1
Expert's answer
2021-07-14T12:02:44-0400

The following table shows the total revenue generated and marginal revenue per trip



a.

The profit maximizing price is $4 since at this point the total revenue is at maximum. The level of output is efficient since the total revenue of 1.6 million is more than the total costs of 1.5 million.

b.

At a price of $4 the company makes a profit of 100,000.

Therefore the company should build the bridge since it will make a profit.

c.

If the government were to build the bridge, they should charge a price equal to to the marginal cost of building the bridge. Hence the price charged should be $0.

d.

If the government charges $0 per trip, then the consumer surplus resulting from the efficient level of production is:

"CS=\frac{1}{2}\times B\times H\\=\frac{1}{2}\times(800,000)\times(\$8-\$0)\\=\$3.2 \space million"


As the value of consumer surplus is greater than the total cost of building the bridge, the government should build the bridge.